OpenStudy (anonymous):

wite a polynomial function with rational coefficients so that p(x)=0 has the given roots 3,-2,5

OpenStudy (lieutenantgeneral):

I will be there in a while but jorgeeeee here's how to expand it (x-5)(x-2i)=0 x(x-2i)-5(x-2i)=0 then use the distributive property to expand it :) let me know if you can't do that too, i'll help you with it :)

OpenStudy (gorv):

(x-3)(x-(-2))(x-5)=0 (x-3)(x+2)(x-5)=0 (x^2-x-6)(x-5)=0 x^3-x^2-6x-5x^2+5x+30=0 x^3-6x^2-x+30=0